# Organize the Web: Constructing Domains Enhances Pre-Training Data Curation

### v2

@@ Abstract
| Website | weborganizer.allen.ai | |
| Artifacts | hf.co/WebOrganizer | |
@@ Appendix A Domain Descriptions
| - Financial crime and litigation fall under ‘Finance & Business’ | |
| - Social issues and the legislative process fall under ‘Politics’ | |
@@
| - Educational pages about a topic, e.g., food or mathematics, fall under that topic | |
@@
| - Music or film discussed as art rather than entertainment falls under ‘Art & Design’ | |
@@
| - Nutritional sciences fall under ‘Health’ | |
@@
| - Health insurance falls under ‘Finance & Business’ | |
@@
| - Includes: raw materials, industrial goods, chemicals, textiles | |
| - General business topics or business finance fall under ‘Finance & Business’ | |
@@
| - Text written in literary style fall under the topic of its contents | |
@@
| - Specific social activity (e.g., sports or board games) fall under those topics | |
@@
| - Detailed descriptions of tourist destinations fall under ‘History’ |
@@
| - Note that the content may appear similar to a ‘Knowledge Article’ in some cases, but is not verified and may contain self-promotion | |
@@
| - Examples: interviews (e.g., in a newspaper), the transcript of a court hearing, movie, podcast, lecture, or speech | |
@@
| - Examples: Community sites like reddit, comment sections on news article or blogs | |
@@
| - Examples: sitemap, product catalog, search results, news listings with short snippets of articles | |
| - Note that hyperlinks are not visible from the text content and have to be deduced | |
@@
| - Unlike ‘Customer Support’, meant for developers and experts, rather than end-users | |
@@
| - Published on a moderated platform (like Wikipedia) or by a reputable source | |
@@
| - Examples: Buzzfeed-style articles, “Top 10” lists, “4 best places to visit in X” | |
| - Lists showing the site contents and facilitate navigation fall under ‘Content Listing’ | |
@@
| - Examples: a press release, a blog post by an organization such as a company, university, government agency, non-profit organization | |
@@
| - Long reads, profiles, editorials, and journalistic essays fall under ‘Nonfiction Writing’ | |
| - Newspaper interviews fall under ‘Audio Transcript’ | |
@@
| - Also includes products in a wider sense, for example university course descriptions | |
@@
| - Examples: a table, datasheet, movie database, glossary, dictionary, json file, csv, xml | |
@@
| - Examples: a troubleshooting guide | |
@@
| - If the page has multiple snippets of truncated articles, choose ’Content Listing’ | |
| - Also includes multimedia web pages where the web page text primarily describes and supplements the audiovisual content, e.g., a video description or image gallery | |
@@
| - The page must contain the actual content of the tutorial / how-to guide | |
| - Guides specific to products/services from the website fall under ‘Customer Support’ | |

### v3

@@ Abstract
Website | weborganizer.allen.ai |
Artifacts | hf.co/WebOrganizer |
@@ Appendix A Domain Descriptions
- Financial crime and litigation fall under ‘Finance & Business’ | |
- Social issues and the legislative process fall under ‘Politics’ | |
@@
- Educational pages about a topic, e.g., food or mathematics, fall under that topic | |
@@
- Music or film discussed as art rather than entertainment falls under ‘Art & Design’ | |
@@
- Nutritional sciences fall under ‘Health’ | |
@@
- Health insurance falls under ‘Finance & Business’ | |
@@
- Includes: raw materials, industrial goods, chemicals, textiles | |
- General business topics or business finance fall under ‘Finance & Business’ | |
@@
- Text written in literary style fall under the topic of its contents | |
@@
- Specific social activity (e.g., sports or board games) fall under those topics | |
@@
- Detailed descriptions of tourist destinations fall under ‘History’ |
@@
- Note that the content may appear similar to a ‘Knowledge Article’ in some cases, but is not verified and may contain self-promotion | |
@@
- Examples: interviews (e.g., in a newspaper), the transcript of a court hearing, movie, podcast, lecture, or speech | |
@@
- Examples: Community sites like reddit, comment sections on news article or blogs | |
@@
- Examples: sitemap, product catalog, search results, news listings with short snippets of articles | |
- Note that hyperlinks are not visible from the text content and have to be deduced | |
@@
- Unlike ‘Customer Support’, meant for developers and experts, rather than end-users | |
@@
- Published on a moderated platform (like Wikipedia) or by a reputable source | |
@@
- Examples: Buzzfeed-style articles, “Top 10” lists, “4 best places to visit in X” | |
- Lists showing the site contents and facilitate navigation fall under ‘Content Listing’ | |
@@
- Examples: a press release, a blog post by an organization such as a company, university, government agency, non-profit organization | |
@@
- Long reads, profiles, editorials, and journalistic essays fall under ‘Nonfiction Writing’ | |
- Newspaper interviews fall under ‘Audio Transcript’ | |
@@
- Also includes products in a wider sense, for example university course descriptions | |
@@
- Examples: a table, datasheet, movie database, glossary, dictionary, json file, csv, xml | |
@@
- Examples: a troubleshooting guide | |
@@
- If the page has multiple snippets of truncated articles, choose ’Content Listing’ | |
- Also includes multimedia web pages where the web page text primarily describes and supplements the audiovisual content, e.g., a video description or image gallery | |
@@
- The page must contain the actual content of the tutorial / how-to guide | |
- Guides specific to products/services from the website fall under ‘Customer Support’ | |